OpenStudy (anonymous):

A population of seals is modeled by the equation Ps= P0 * 5^t/3, where t is measured in years. Find a. The time required for the population to double in size b. the percentage increase of the population during the first four years

OpenStudy (anonymous):

You have.\[P_s(t)=P_0 \cdot 5^{t/3}\]To do part (a), you need to know the original population (hint, evaluate at \(t=0\)). Then, double it, and set \(P_s(t)\) to that doubled value, and find the \(t\) value that satisfies that. For part (b), percentage increase is given as follows.\[\text{Percent Increase}=\left(\frac{\text{new}-\text{original}}{\text{original}}\right)\cdot 100\%\]Plug in the original value where it says original, and \(P_s(t)\) where it says new. Does that make sense?

OpenStudy (anonymous):

Yes thank you.